# Quantifying Resolution Limits in Pedestal Profile Measurements with Gaussian Process Regression

### v1

@@ Abstract
Edge transport barriers (ETBs) in magnetically confined fusion plasmas, commonly known as pedestals, play a crucial role in achieving high confinement plasmas. However, their defining characteristic, a steep rise in plasma pressure over short length scales, makes them challenging to diagnose experimentally. In this work, we use Gaussian Process Regression (GPR) to develop first-principles metrics for quantifying the spatiotemporal resolution limits of inferring differentiable profiles of temperature, pressure, or other quantities from experimental measurements. Although we focus on pedestals, the methods are fully general and can be applied to any setting involving the inference of profiles from discrete measurements. First, we establish a correspondence between GPR and low-pass filtering, giving an explicit expression for the effective ‘cutoff frequency’ associated with smoothing incurred by GPR. Second, we introduce a novel information-theoretic metric, $N_{eff}$, which measures the effective number of data points contributing to the inferred value of a profile or its derivative. These metrics enable a quantitative assessment of the trade-off between ‘over-fitting’ and ‘over-regularization’, providing both practitioners and consumers of GPR with a systematic way to evaluate the credibility of inferred profiles. We apply these tools to develop practical advice for using GPR in both time-independent and time-dependent settings, and demonstrate their usage on inferring pedestal profiles using measurements from the DIII-D tokamak.
  • July 3, 2025
@@ 1 Introduction
A major goal in magnetic confinement fusion energy is the development of predictive models of transport for energy, particles, and momentum across transport barriers. Edge transport barriers (ETBs), generally known as pedestals, are of particular interest due to their critical role in the performance of H-mode and other enhanced confinement regimes [1, 2]. Transport across pedestals must balance good confinement in the plasma core with a multitude of competing requirements such as heat flux management to the divertor and impurity accumulation, constituting the so-called “core-edge integration” challenge [3, 4, 5]. Although ETBs are the primary focus of this work, internal transport barriers (ITBs) have also been explored as ways to improve the performance of fusion devices and meet these challenges [6], and the tools developed here are applicable to them as well.
@@
Achieving reliable control and quantification of trade-offs of transport in pedestals and other transport barriers requires experimentally validated models of physical processes in these regions [7, 8, 9]. However, the very property which makes transport barriers attractive for confinement, their steep rise in plasma pressure over a short length scale, makes them challenging to diagnose experimentally. The typical radial width of a pedestal $\ell$ is only a small fraction of the device size $L$, $\ell/L\sim 2-10\%$. For diagnostics such as Thomson scattering, only a few channels might cover the width of the pedestal, limiting the accuracy to which profile gradients can be inferred from experimental measurements. Meanwhile, transport mechanisms frequently exhibit critical gradient phenomena where turbulent fluxes become extremely sensitive to gradients of density, temperature, or pressure above certain instability thresholds. Transport mechanisms are also sensitive to velocity shear, which can act to stabilize or entirely quench transport. In cases where the radial electric field is dominated by diamagnetic flow $E_{r}\sim\nabla p$, computing $E\times B$ shear requires accurately inferring the second derivative of pressure profiles from measurements.
@@
Further complicating matters in pedestal modeling is the interest in understanding temporally-dependent phenomena such as the edge-localized mode (ELM) cycle [10, 11] or the L-H transition [12], which can involve fast temporal evolution of pedestal profiles. These time-dependent phenomena relate to problems such as the maximum pedestal pressure which can be achieved after an ELM, as well the physics of the minimum power threshold necessary for achieving H-mode. Beyond pedestals, there is growing interest in direct profile measurements of mesoscale phenomena, such as ‘staircase’ profiles in L-mode and H-mode [13, 14, 15, 16], as well as profile variation due to magnetic islands [17, 18].
@@
In this work we address the problem of uncertainty quantification in inferences of differentiable profiles of plasma properties from spatially localized measurements. The key question we address is: what is the maximum spatial and temporal scale of features which can be reliably inferred from given a set of profile measurements? To make this problem quantitative, we adopt the popular approach of Gaussian Process Regression (GPR) [19, 20, 21, 22, 23]. GPR is enormously flexible, being used to perform principled inference and uncertainty quantification from complex experimental setups [24, 25, 26, 27], produce tomographic inversions [28, 29, 30, 31, 32], and provide surrogate models of complex phenomena [33, 34].
@@
GPR has the important advantages that it does not require a functional form of the fit to be specified ahead of time, and can also provide statistically principled error quantification for derivatives. Although GPR is often called ‘non-parameteric’, we discuss how one of the main challenges of interpreting results from GPR is the problem of kernel and kernel hyperparmeter selection. In particular, we discuss how pushing the resolution limits using GPR necessarily involves a trade-off between ‘over-regularization’ and ‘over-fitting’.
@@
Rather than identify a single best choice of kernel, $N_{eff}$ provides an interpretable information-theoretic metric to detect overfitting and compare the credibility of different fits. Together, these two results provide tools to replace trial-and-error in the fitting process with quantitative metrics that can be used to engineer the tradeoff between over-fitting and over-regularization – something that simpler parametric fits cannot provide.
This paper is organized as follows: We begin in section 2 by introducing notation and conventions for Gaussian Process Regression adopted in this work. Afterwards, readers who are interested in applications of the results can skip ahead to section 5, where we demonstrate their usage in a practical situation involving pedestal profiles on DIII-D. Otherwise, we follow in section 3 with a discussion of how GPR can be viewed as a form of low-pass filtering and demonstrate the insufficiency of the kernel length scale in determining the length scale associated with this filtering. In section 4 we derive an expression for $N_{eff}$ and provide several main properties it satisfies. In section 5 we demonstrate the usage of these tools in practical situations demonstrating how they affect profile fits. Finally, in section 6 we summarize the work and discuss further potential applications enabled by the results.
@@ 2 Profile Fitting with GPR
In this section, we introduce some background on GPR and discuss how the choice of GPR methodology can influence the resulting profiles inferred from the measurements. In section 2.1 we fix notation and give a conceptual overview of Gaussian Processes and the relation to GPR. Then in section 2.2 we introduce the problem of hyperparameter selection in GPR.
@@ 2.1 Conceptual Overview of GPR
In general, functions have an infinite number of degrees of freedom, whereas the number of measurements we have is decidedly finite. Hence, the problem of inferring $f(x)$ from measurements is in some sense always severely underdetermined, so we must always make some choice about how to ‘fill in the gaps’ between measurements. Gaussian Process Regression (GPR), also known as kriging in the literature, is a particular choice of how to resolve this severe underdetermination which has the nice properties of (1) not constraining the functional form of the fit, and (2) having a principled way to compute statistical uncertainties for the resulting function $f(x)$ and any of its derivatives [19].
Several works give extensive discussions of GPR and how to apply it in the context of profile fitting [20, 23], so here we will give a short overview of GPR and the key results we will use. Conceptually, a Gaussian process (GP) is a way of assigning a probability measure to some ensemble (i.e. set) of functions $H$, with each realization $f\in H$ being a different possible function $f(x)$. Similar to how the distribution of a Gaussian random vector is entirely specified by the mean and covariance matrix of its components, a GP is entirely specified by a mean $\mu(x)$ and covariance kernel $k(x,x^{\prime})$, which satisfy
@@
where $\boldsymbol{\beta}(x)=[\beta_{1}(x),...,\beta_{N}(x)]^{T}$ is the vector of so-called ‘regression coefficients’. The posterior covariance, defined as the conditional covariance $\hat{k}(x,x^{\prime}):=\operatorname{Cov}[f(x),f(x^{\prime})|\mathbf{Y}]$, also has an explicit expression
@@
For concreteness, for the prior we will focus on the family of stationary Matérn kernels [19]. In 1D, these kernels are given by $k(x,x^{\prime})=\kappa(x-x^{\prime})$ with
@@ 2.2 Hyperparameter Selection
In this section, we first provide some intuition for the Matérn kernel hyperparameters. We follow this with a brief overview of different methods for choosing or marginalizing over hyperparmeters. In this work we remain agnostic to how the hyperparameters are chosen or marginalized over, as the methods developed in this work focus on how to interpret any set of hyperparameters from any method. For more detailed prescriptions for hyperparameter inference, readers are directed to the references [20, 23].
The easiest hyperparameter to understand is the variance parameter $\sigma^{2}$, which determines the typical amplitude of $f(x)$ via the relation
@@
It is also clear from this that $\sigma^{2}$ has the same dimensions as $f^{2}$. Next, the smoothness hyperparameter $\nu$, which is dimensionless, controls the differentiability of the realizations $f\in H$. In general $f$ will be $\lceil\nu\rceil-1$ times differentiable in the mean-square sense [19]. Finally, the length scale parameter $\ell$ has the same dimensions as the measurement variable $x$.
@@
Generally speaking, for fixed values of $\sigma^{2}$ and $\nu$, as the length scale $\ell$ is decreased, the inferred profile picks up on smaller-scale fluctuations present in the data $\mathbf{Y}$. An example of this is shown in figure 1, which illustrates how two example inferred profiles using the same data are decomposed into regression coefficients $\beta_{i}(x)$ for different values of $\ell$. Notice that for the case with larger $\ell$, the functions $\beta_{i}(x)$ are broader, which smooth out fluctuations in the data $\mathbf{Y}$. Meanwhile, for the case with smaller $\ell$, the $\beta_{i}(x)$ are narrower, which lead to larger fluctuations in the inferred profile.
@@
Figure 1: (Top) Illustration of how the length scale $\ell$ parameter affects profiles (orange curve with ±$\pm 34\%$ confidence intervals) inferred using GPR. The decomposition of this profile into scaled regression coefficient functions $Y_{i}\beta_{i}(x)$ (smooth curves, various colors) associated with the data points $Y_{i}$ (scattered points, various colors) is also shown. (Bottom) Plot of the $N_{eff}(x)$ metric defined in section 4. The colored vertical lines show the $x$ location of the data points. $N_{eff}\approx 1$ indicates only one measurement is contributing to the inference of the profile at a given point, indicating overfitting in the $\ell=0.12$ case.
@@
Given the influence of the hyperparameters on the inferred profile in GPR, care must be taken in how to specify the hyperparameters. The error bars shown in figure 1 are rigorous and exact within the assumptions of the uncertainty model. However, any calculation is only as good as its model! A fully statistically principled approach requires a more careful examination of the assumptions underlying the inference process.
@@
Broadly speaking, one can attempt to either consider some specific choice of hyperparameter $\boldsymbol{\theta}^{*}$, or one can consider a distribution of hyperparameters $p_{\boldsymbol{\theta}}$ and consider an ensemble of fits $\hat{f}_{\boldsymbol{\theta}}(x)=\mathbb{E}[f(x)|\boldsymbol{\theta},\mathbf{Y}]$ with different values of $\boldsymbol{\theta}$. In the first case, the hyperparameter might be chosen to optimize some objective function, such as a maximum likelihood estimator (MLE) which tries to maximize the likelihood $p(\mathbf{Y};\boldsymbol{\theta})$ of observing the data $\mathbf{Y}$ with prior GP specified by $\boldsymbol{\theta}$,
$\boldsymbol{\theta}_{MLE}:=\operatorname{argmax}_{\boldsymbol{\theta}}p(\mathbf{Y};\boldsymbol{\theta})$
@@
In the latter case, the inferred profile is usually taken to be the average over the ensemble of profiles $\hat{f}(x)=\mathbb{E}[\hat{f}_{\boldsymbol{\theta}}(x)|\mathbf{Y}]$, which is known as marginalization.
Previous works have done comparisons between different methodologies for specifying $\boldsymbol{\theta}$, in particular optimization versus marginalization [20]. In the following however, we will focus on the quality of a fit for a given $\boldsymbol{\theta}$ independent of how the hyperparameters are chosen. In the case of an optimization method or similar providing a single ‘best’ choice of hyperparameter, then the quality of the fit can be directly evaluated, potentially suggesting a need to update bounds on the optimization or assumptions on the hyperprior. In the case of marginalization, these metrics will allow the quality of fits within the ensemble to be evaluated, allowing different fits in the ensemble to be weighted or rejected/accepted based on this quality metric, and also guide the design of prior distributions for the hyperparameters.

### v2

@@ Abstract
Edge transport barriers (ETBs) in magnetically confined fusion plasmas, commonly known as pedestals, play a crucial role in achieving high confinement plasmas. However, their defining characteristic, a steep rise in plasma pressure over short length scales, makes them challenging to diagnose experimentally. In this work, we use Gaussian Process Regression (GPR) to develop first-principles metrics for quantifying the spatiotemporal resolution limits of inferring differentiable profiles of temperature, pressure, or other quantities from experimental measurements. Although we focus on pedestals, the methods are fully general and can be applied to any setting involving the inference of profiles from discrete measurements. First, we establish a correspondence between GPR and low-pass filtering, giving an explicit expression for the effective ‘cutoff frequency’ associated with smoothing incurred by GPR. Second, we introduce a novel information-theoretic metric, $N_{eff}$, which measures the effective number of data points contributing to the inferred value of a profile or its derivative. These metrics enable a quantitative assessment of the trade-off between ‘over-fitting’ and ‘over-regularization’, providing both practitioners and consumers of GPR with a systematic way to evaluate the credibility of inferred profiles with respect to their degree of regularization. We apply these tools to develop practical advice for using GPR in both time-independent and time-dependent settings, and demonstrate their usage on inferring pedestal profiles using measurements from the DIII-D tokamak.
  • December 18, 2025
@@ 1 Introduction
A major goal in magnetic confinement fusion energy is the development of predictive models of transport for energy, particles, and momentum across transport barriers. Edge transport barriers (ETBs), generally known as pedestals, are of particular interest due to their critical role in the performance of H-mode and other enhanced confinement regimes [wagner_quarter-century_2007, groebner_elements_2023]. Transport across pedestals must balance good confinement in the plasma core with a multitude of competing requirements such as heat flux management to the divertor and impurity accumulation, constituting the so-called “core-edge integration” challenge [dux_accumulation_2003, pitts_physics_2019, viezzer_prospects_2023]. Although ETBs are the primary focus of this work, internal transport barriers (ITBs) have also been explored as ways to improve the performance of fusion devices and meet these challenges [ida_internal_2018], and the tools developed here are applicable to them as well.
@@
Achieving reliable control and quantification of trade-offs of transport in pedestals and other transport barriers requires experimentally validated models of physical processes in these regions [terry_validation_2008, greenwald_verification_2010, fischer_estimation_2020]. However, the very property which makes transport barriers attractive for confinement, their steep rise in plasma pressure over a short length scale, makes them challenging to diagnose experimentally. The typical radial width of a pedestal $\ell$ is only a small fraction of the device size $L$, with $\ell/L\sim 2\text{--}10\%$. For diagnostics such as Thomson scattering, only a few channels might cover the width of the pedestal, limiting the accuracy to which profile gradients can be inferred from experimental measurements. Meanwhile, transport mechanisms frequently exhibit critical gradient phenomena where turbulent fluxes become extremely sensitive to gradients of density, temperature, or pressure above certain instability thresholds. Transport mechanisms are also sensitive to velocity shear, which can act to stabilize or entirely quench transport. In cases where the radial electric field is dominated by diamagnetic flow $E_{r}\sim\nabla p$, computing $E\times B$ shear requires accurately inferring the second derivative of pressure profiles from measurements.
@@
Further complicating matters in pedestal modeling is the interest in understanding temporally-dependent phenomena such as the edge-localized mode (ELM) cycle [laggner_inter-elm_2019, diallo_review_2020] or the L-H transition [schmitz_role_2012], which can involve fast temporal evolution of pedestal profiles. These time-dependent phenomena relate to problems such as the maximum pedestal pressure which can be achieved after an ELM, as well the physics of the minimum power threshold necessary for achieving H-mode. Beyond pedestals, there is growing interest in direct profile measurements of mesoscale phenomena, such as ‘staircase’ profiles in L-mode and H-mode [dif-pradalier_finding_2015, ashourvan_formation_2019, liu_evidence_2021, choi_mesoscopic_2024], as well as profile variation due to magnetic islands [waelbroeck_theory_2009, choi_interaction_2021].
@@
In this work we address the problem of uncertainty quantification in inferences of differentiable profiles of plasma properties from spatially localized measurements. The key question we address is: what is the maximum spatial and temporal scale of features which can be reliably inferred from a given set of profile measurements? To make this problem quantitative, we adopt the popular approach of Gaussian Process Regression (GPR) [rasmussen_gaussian_2008, chilenski_improved_2015, ho_application_2019, pavone_machine_2023, michoski_gaussian_2024]. GPR is enormously flexible, being used to perform principled inference and uncertainty quantification from complex experimental setups [chilenski_experimentally_2017, cao_hysteresis_2019, nishizawa_plasma_2021, kwak_bayesian_2022], produce tomographic inversions [li_bayesian_2013, langenberg_forward_2016, wang_bayesian_2019, chao_gaussian_2020, matos_deep_2020], and provide surrogate models of complex phenomena [preuss_gaussian_2016, rodriguez-fernandez_nonlinear_2022].
@@
GPR has two important advantages: it does not impose a fixed functional form for the prior distribution of functions which can be inferred, and it can provide statistically principled error quantification for both the inferred profiles and their derivatives. This can be contrasted with ‘parametric’ regression techniques, such as spline fits or the modified hyperbolic tangent functions commonly used in pedestal modeling [groebner_scaling_1998, groebner_progress_2001], which constrain the inferred profiles to lie within a predetermined family of functions characterized by a finite number of parameters. While GPR is often called ‘non-parameteric’, we discuss how one of the main challenges of interpreting results from GPR is the problem of kernel and kernel hyperparmeter selection, as these choices often make significant impacts on the resulting inferred profiles. In particular, we discuss how pushing the resolution limits using GPR necessarily involves a trade-off between ‘over-regularization’ and ‘over-fitting’.
@@
Rather than identify a single best choice of kernel, $N_{eff}$ provides an interpretable information-theoretic metric useful for detecting when a given kernel is vulnerable to over-fitting. Together, these two results provide tools to replace trial-and-error during the fitting process or hyperprior development process with quantitative metrics that can be used to engineer the tradeoff between over-fitting and over-regularization – something that simpler parametric fits cannot provide.
This paper is organized as follows: We begin in section 2 by introducing notation and conventions for Gaussian Process Regression adopted in this work, and discuss trade-offs in GPR methodology. Afterwards, readers who are interested in applications of the results can skip ahead to section 5, where we demonstrate their usage in a practical situation involving pedestal profiles on DIII-D. Otherwise, we follow in section 3 with a discussion of how GPR can be viewed as a form of low-pass filtering and demonstrate the insufficiency of the kernel length scale in determining the length scale associated with this filtering. In section 4 we derive an expression for $N_{eff}$ and provide several main properties it satisfies. In section 5 we demonstrate the usage of these tools in practical situations demonstrating how they affect profile fits. Finally, in section 6 we summarize the work and discuss further potential applications enabled by the results.
@@ 2 Profile Fitting with GPR
In this section, we introduce some background on GPR and discuss how the choice of GPR methodology can influence the resulting profiles inferred from the measurements. In section 2.1 we fix notation and give a conceptual overview of Gaussian Processes and the relation to GPR. Then in section 2.2 we introduce the problem of hyperparameter selection in GPR, in particular focusing on the trade-off between over-regularization and over-fitting inherent to all profile fitting techniques.
@@ 2.1 Conceptual Overview of GPR
In general, functions have an infinite number of degrees of freedom, whereas the number of measurements we have is decidedly finite. Hence, the problem of inferring $f(x)$ from measurements is in some sense always severely underdetermined, so we must always make some choice about how to ‘fill in the gaps’ between measurements. Gaussian Process Regression (GPR), also known as kriging in the literature, is a particular choice of how to resolve this severe underdetermination which has the nice properties of (1) not imposing a fixed functional form for the prior distribution of functions which can be inferred, and (2) having a principled way to compute statistical uncertainties for the resulting function $f(x)$ and any of its derivatives [rasmussen_gaussian_2008].
Several works give extensive discussions of GPR and how to apply it in the context of profile fitting [chilenski_improved_2015, michoski_gaussian_2024], so here we will give a short overview of GPR and the key results we will use. Conceptually, a Gaussian process (GP) is a way of assigning a probability measure to some ensemble (i.e. set) of functions $H$, with each realization $f\in H$ being a different possible function $f(x)$. Similar to how the distribution of a Gaussian random vector is entirely specified by the mean and covariance matrix of its components, a GP is entirely specified by a mean $\mu(x)$ and covariance kernel $k(x,x^{\prime})$, which satisfy
@@
Typically the ensemble $H$ can be extended to include all functions with square-integrable derivatives up to some given order, i.e. a Sobolev space.
@@
where $\boldsymbol{\beta}(x)=[\beta_{1}(x),...,\beta_{N}(x)]^{T}$ is the vector of so-called ‘regression coefficients’. Observe that the number of regression coefficients grows as the number of data points grows, which allows the inferred profile to become more flexible as more data is added. The posterior covariance, defined as the conditional covariance $\hat{k}(x,x^{\prime}):=\operatorname{Cov}[f(x),f(x^{\prime})|\mathbf{Y}]$, also has an explicit expression
@@
For concreteness, for the prior we will focus on the family of stationary Matérn kernels [rasmussen_gaussian_2008]. In 1D, these kernels are given by $k(x,x^{\prime})=\kappa(x-x^{\prime})$ with
@@ 2.2 Hyperparameter Selection
In this section, we outline how hyperparameters affect the behavior of GPR inferences and discuss potential issues that can arise if they are not properly chosen. We begin by providing some intuition for the Matérn kernel hyperparameters. Then, we summarize several common methods for choosing or marginalizing over hyperparmeters, and discuss differences in the uncertainty estimates and interpretations of inferences from these methods. Finally, we discuss the need for quantitative metrics to assess whether a chosen set of hyperparameters lead to over-regularization or over-fitting, giving criteria to determine if an inference represents a credible interpretation of the data. In this work we remain agnostic to how the hyperparameters are chosen or marginalized over, as the methods developed in this work focus on how to interpret any set of hyperparameters from any method. For more detailed prescriptions for hyperparameter inference, readers are directed to the references [chilenski_improved_2015, michoski_gaussian_2024].
Recall that the kernel $k$ determines the prior distribution of functions $f$ over which the GPR inference is carried out. Thus, different choices of hyperparameter encode different assumptions about what the ‘typical’ $f$ looks like. The easiest hyperparameter to understand is the variance parameter $\sigma^{2}$, which determines the typical amplitude of $f(x)$ via the relation
@@
It is also clear from this that $\sigma^{2}$ has the same dimensions as $f^{2}$. Next, the smoothness hyperparameter $\nu$, which is dimensionless, controls the differentiability of the realizations $f\in H$. In general $f$ will be $\lceil\nu\rceil-1$ times differentiable in the mean-square sense [rasmussen_gaussian_2008]. Finally, the length scale parameter $\ell$ has the same dimensions as the measurement variable $x$.
@@
Generally speaking, for fixed values of $\sigma^{2}$ and $\nu$, as the length scale $\ell$ is decreased, the inferred profile picks up on smaller-scale fluctuations present in the data $\mathbf{Y}$. An example of this is shown in figure 1, which illustrates how two example inferred profiles using the same data are decomposed into regression coefficients $\beta_{i}(x)$ for different values of $\ell$. Notice that for the case with larger $\ell$, the functions $\beta_{i}(x)$ are broader, which smooth out fluctuations in the data $\mathbf{Y}$. Meanwhile, for the case with smaller $\ell$, the $\beta_{i}(x)$ are narrower, which lead to larger fluctuations in the inferred profile. As we will see later in section 3, this difference in behavior explicitly reflects the bias that the assumptions on $f$ introduce into the GPR estimate $\hat{f}$.
@@
Figure 1: (Top) Illustration of how the length scale $\ell$ parameter affects profiles (orange curve with $\pm 34\%$ confidence intervals) inferred using GPR. The decomposition of this profile into scaled regression coefficient functions $Y_{i}\beta_{i}(x)$ (smooth curves, various colors) associated with the data points $Y_{i}$ (scattered points, various colors) is also shown. (Bottom) Plot of the $N_{eff}(x)$ metric defined in section 4. The colored vertical lines show the $x$ location of the data points. $N_{eff}\approx 1$ indicates only one measurement is contributing to the inference of the profile at a given point, indicating overfitting in the $\ell=0.12$ case.
@@
Given the influence of the hyperparameters on the inferred profile in GPR, care must be taken in how to specify the hyperparameters. The error bars shown in figure 1 are rigorous and exact within the assumptions of the uncertainty model. However, any calculation is only as good as its model! A fully statistically principled approach requires a more careful examination of the assumptions underlying the inference process. Here, we first outline the approaches by which hyperparameters are usually determined, then discuss necessary conditions for the physical credibility of the inferences produced by these approaches.
@@
Broadly speaking, one might either attempt to find a single ‘best’ choice of hyperparameter $\boldsymbol{\theta}^{*}$, or one might consider a distribution of hyperparameters $p_{\boldsymbol{\theta}}$ and consider an ensemble of fits $\hat{f}_{\boldsymbol{\theta}}(x)=\mathbb{E}[f(x)|\boldsymbol{\theta},\mathbf{Y}]$ with different values of $\boldsymbol{\theta}$. In the first case, the hyperparameter might be chosen to optimize some objective function, such as a maximum likelihood estimator (MLE)
$\boldsymbol{\theta}_{MLE}:=\operatorname{argmax}_{\boldsymbol{\theta}}p(\mathbf{Y}|\boldsymbol{\theta})$
which tries to maximize the likelihood $p(\mathbf{Y}|\boldsymbol{\theta})$ of observing the data $\mathbf{Y}$ with prior GP specified by a given value $\boldsymbol{\theta}$ of the hyperparameters. $p(\mathbf{Y}|\boldsymbol{\theta})$ is also known as the marginal likelihood, since it involves marginalizing over realizations $f$, and usually the optimization is carried out over the log marginal likelihood for numerical reasons.
In the latter case, typically a hyperprior $q_{\boldsymbol{\theta}}$ is first specified, then Bayes’ theorem is applied to find a posterior distribution which weights the hyperprior by the likelihood of observing the data
$p_{\boldsymbol{\theta}}(\boldsymbol{\theta}|\mathbf{Y})=q_{\boldsymbol{\theta}}(\boldsymbol{\theta})\frac{p(\mathbf{Y}|\boldsymbol{\theta})}{\int p(\mathbf{Y}|\boldsymbol{\theta})q(\boldsymbol{\theta})d\boldsymbol{\theta}}$
@@
The inferred profile is then taken to be the average over the ensemble $p_{\boldsymbol{\theta}}$ of profiles $\hat{f}(x)=\mathbb{E}[\hat{f}_{\boldsymbol{\theta}}(x)|\mathbf{Y}]$, an approach known as (Bayesian) marginalization.
Previous works have done comparisons between different methodologies for specifying $\boldsymbol{\theta}$, in particular optimization versus marginalization [chilenski_improved_2015]. The difference between the uncertainty estimates produced by the two approaches can be understood through the law of total variance: If $\sigma^{2}_{\boldsymbol{\theta}}(x):=\operatorname{Var}[f(x)|\boldsymbol{\theta},\mathbf{Y}]$ is the uncertainty of an inference given a single point estimate of $\boldsymbol{\theta}$, then the total uncertainty can be computed
$\operatorname{Var}[f(x)|\mathbf{Y}]=\mathbb{E}[\sigma^{2}_{\boldsymbol{\theta}}(x)|\mathbf{Y}]+\operatorname{Var}[\hat{f}_{\boldsymbol{\theta}}(x)|\mathbf{Y}]$ (5)
where the variance and expectation are taken over $p_{\boldsymbol{\theta}}$.
The first term $\mathbb{E}[\sigma^{2}_{\boldsymbol{\theta}}(x)|\mathbf{Y}]$ on the right hand side of (5) is sometimes referred to as aleatoric (within-model) uncertainty, and captures the averaged effect of how measurement uncertainties modeled by $\boldsymbol{\varepsilon}$ propagate through to uncertainty in the inference for a given choice of kernel and associated hyperparameters $\boldsymbol{\theta}$. Note that GPR with a specified kernel is already a Bayesian method which posits an explicit hypothesis for the prior distribution of profiles $f$ over which the inference is performed, hence it is appropriate to describe its output uncertainty as aleatoric.
The second term $\operatorname{Var}[\hat{f}_{\boldsymbol{\theta}}(x)|\mathbf{Y}]$ is referred to as epistemic (between-model) uncertainty, and captures uncertainty arising from variations in the ensemble of inferences due to varying choices of $\boldsymbol{\theta}$. If the variation of $\boldsymbol{\theta}$ within the distribution $p_{\boldsymbol{\theta}}$ produces variations in the inference $\hat{f}_{\boldsymbol{\theta}}(x)$ that are larger than the typical uncertainty $\sigma^{2}_{\boldsymbol{\theta}}(x)$ for fixed $\boldsymbol{\theta}$, then the epistemic uncertainty will dominate the aleatoric uncertainty.
In the remainder of this work, we will primarily focus on GPR inferences for individual values of $\boldsymbol{\theta}$. These values might arise as the output of an optimization routine, or could represent individual members of an ensemble $p_{\boldsymbol{\theta}}$. When we present error bars for a fixed value of $\boldsymbol{\theta}$, they should be understood as only representing the aleatoric component of the uncertainty. Typically the marginalization approach will result in larger uncertainties due to the inclusion of epistemic uncertainty, especially for gradients.
As discussed earlier, profile fitting is severely undetermined for any finite number of measurements. Typically, this problem must be addressed by including a degree of either implicit or explicit regularization during the profile inference process. When adding regularization, two main issues must be avoided:
  1. 1.
    (Over-regularization) The inference procedure should not inadvertently filter out structures or features present over the spatial or temporal scales of interest.
  2. 2.
    (Over-fitting) The inference procedure needs to ensure that a sufficient diversity of measurements contribute to the inference of the profile at any given point, otherwise the fit may become overly sensitive to noise or systematic error in the measurements.
Avoiding these issues is a necessary (although perhaps not sufficient) condition for the inferred profile to represent a credible interpretation of the measurements.
Our aim is to develop criteria for evaluating whether an inference for a given value of $\boldsymbol{\theta}$ suffers from either over-regularization or over-fitting. This involves translating the mathematical modeling choices made in GPR into an interpretable set of metrics which can be used to assess the credibility of an inference by comparing its degree of regularization with physically meaningful quantities. Intuitively, different values of $\boldsymbol{\theta}$ lead to a tradeoff between over-regularization and over-fitting. It may be the case that there are no admissible values of the hyperparameter which can simultaneously meet the requirements to avoid over-regularization and over-fitting, in which case either more measurements or less demanding resolution requirements are needed.
In the context of optimization or point estimates of $\boldsymbol{\theta}$, these metrics can inform the bounds on admissible values of the hyperparameter and determine if an individual inference can be taken as a credible (although perhaps non-exhaustive) interpretation of the data. We will discuss these applications in more detail in the following sections. Although we do not study marginalization in detail, these metrics can also be used to guide the design and interpretation of hyperpriors on $\boldsymbol{\theta}$. With a properly designed hyperprior, the fraction of hyperparameters in the posterior ensemble $p_{\boldsymbol{\theta}}$ which meet the credibility criteria can be used to give an overall evaluation of the credibility of the inference.